# Pruning One More Token is Enough: Leveraging Latency-Workload Non-Linearities for Vision Transformers on the Edge

### v2

@@ 3.3 Token Pruning Method
The schedule selection described in Secs. 3.1 and 3.2 identifies the location and number of tokens to prune. Now we give a training-free token pruning mechanism to decide which tokens to prune. This pruning mechanism is applied during inference (Fig. 3). The pruning schedule, computed offline, is an input to our token pruning mechanism.
In terms of design goals, we want our pruning mechanism to require (1) no finetuning on $M$ and (2) to introduce minimal overhead. First, requiring no fine-tuning or training is a convenience for practitioners, and allows pre-trained models to be deployed readily. Second, keeping overhead to a minimum for small workload sizes is necessary to yield latency reductions as previously mentioned in 3.2.
The primary design choice of pruning techniques is choosing a method to rank the importance of tokens. Similarly to other pruning techniques [3, 15], we rank the importance of each token by measuring the quantity of attention received from all other tokens.
Figure 5: Illustration of our token pruning method for a single transformer encoder layer at inference time. We compute an attention score and V score for each token to estimate importance, then prune the least informative tokens. Pruned tokens are aggregated into a single ‘garbage bin’ or inattentive token to retain information content as in EViT [19].
Our method incorporates two aspects of importance, one from the $A_{\text{H}}$ matrix and one from the $V$ matrix. For the $A_{\text{H}}$ term, we proceed as follows. Recall that the core attention mechanism of transformers [40] is parameterized by the $Q$, $K$, and $V$ matrices. The attention matrix (including softmax) of a transformer encoder layer with $N$ tokens and $H$ heads is given by $A_{\text{H}}\in\mathbb{R}^{H\times N\times N}$  [40]. To convert the softmax attention matrix $A_{\text{H}}$ into tokenwise scores we: (1) reduce the head dimension $H$ with amax [35] (2) sum along the row $N$ dimension (3) normalize scores by dividing by the max. This transforms $A_{\text{H}}$ into tokenwise scores $\text{A}_{\text{S}}\in\mathbb{R}^{N}$, similar to a component of EViT’s pruning mechanism [19] or Top-K [13].
We then consider a second importance term based on the $V\in\mathbb{R}^{H\times N\times C}$ matrix. This term assigns higher scores to tokens who have high weighting in the $V$ matrix, which marginally increased accuracy. We perform a similar operation for this term: (1) reduce the $H$ dimension of $V$ with amax [35] (2) sum along the feature dimension $C$ (3) normalize scores with softmax. The resulting score vector is $V_{\text{S}}\in\mathbb{R}^{N}$, which we add to $\text{A}_{\text{S}}$ for our overall score:
Token Importance $=\text{A}_{\text{S}}+\text{V}_{\text{S}}.$ (4)
Finally, we borrow from EViT [19] and instead of discarding the pruned tokens, we create a new “inattentive” token based on the averaged features of all pruned tokens and append it to the set of kept tokens. This allows information from pruned tokens to be preserved in this inattentive token, which EViT found improved accuracy. Fig. 5 illustrates the token pruning process for a ViT encoder block or layer.
@@ 4 Evaluation
In this section, we demonstrate the effectiveness of our token pruning schedule and pruning mechanism across devices (listed in Tab. 3) and across various workload sizes. We primarily focus on smaller workloads, since these are representative of edge deployment use-cases [18]. To justify our design decisions, we investigate our $\alpha$ hyperparameter and pruning location choice for our schedule selection and pruning mechanism in Sec. 4.2 and Sec. 4.3, respectively. Then, we compare with a variety of existing token removal methods such as DynamicViT [36], ToMe [1]. We use Top-K [19, 13] as a baseline on ImageNet1K [9]. Sec. 4.4 demonstrates how our method advances the accuracy vs. latency pareto front across devices, workload sizes, and models. Finally, we identify some pitfalls of existing token pruning techniques for small workload sizes in Sec. 4.5. Particularly we show that some existing methods actually increase latency rather than reduce it.
@@
4.1 Experimental Setup
Evaluation is done using three devices with varying hardware characteristics (Tab. 3). We use NVIDIA TX2 [29] and NVIDIA AGX Orin [24] development boards which are edge devices designed for machine learning. Additionally, evaluation is done on a server-grade system with an NVIDIA A100 GPU [25] to demonstrate our method can generalize across a wide range of hardware, even though we design our method for edge devices. Experiments performed on the TX2 and Orin were performed with fixed CPU and GPU clock rates for reproducibility and consistency. The A100 system clocks were unable to be locked because the servers machines are university-wide resources.
In order to measure latency for evaluation, we use the PyTorch benchmark module [33]. All accuracy measurements were performed on the A100 system with batch size 512 on the evaluation split of ImageNet1K. In subsequent experiments, we decide the pruning hyperparameters of each token pruning method for fair comparison with our method. For reproducibility, our code is open source.
4.2 Pruning schedule ablation study across $\alpha$ (utility weight of latency/accuracy)
In 3.1 we introduce an algorithm to decide a number of tokens to prune $R$ according to the GPU tail effect.
Device $\alpha$ R Acc. Median
@@ 4.2 Pruning schedule ablation study across $\alpha$ (utility weight of latency/accuracy)
The hyperparameter of this algorithm is $\alpha$, which determines the relative weighting of the accuracy and latency utilities. Setting to $\alpha=1$ prioritizes accuracy, and only a few tokens might be pruned. Setting to $\alpha=0$ prioritizes latency, and would prune all or nearly-all tokens. To understand the performance spectrum, we performed an ablation study. The data is in Tab. 4.
Evaluation on the AGX Orin for $\alpha>0.5$ yielded $R=193$ (over 75% of tokens pruned), which degraded accuracy by $\sim$2.8for 2.1% less latency. We consider this to be an unfavorable tradeoff. This ablation, in addition to results in Sec. 4.4, suggest that $\alpha\leq 0.5$ is a good choice. Thus we use $\alpha=0.5$ for all evaluations that appear in this work. Intuitively speaking this means accuracy degradation and latency reduction are considered with equal weight according to Eq. 3 when selecting $R$.
4.3 Pruning location ablation study
In 3.2 we decide at which layer our pruning mechanism should be applied. To provide insight into the potential pruning locations, we performed an ablation study. Tab. 5 illustrates latency and accuracy tradeoffs for various pruning locations of DinoV2-G.
Batch Pruning Acc. Median
Size Layer Loss Latency (ms)
1     01/40 3.19 68.4
10/40 1.07 81.3
20/40 0.58 93.4
30/40 0.49 104.4
2     01/40 7.08 79.1
10/40 2.04 104.7
20/40 0.97 133.0
30/40 0.83 160.8
Table 5: Latency/accuracy tradeoffs by pruning location. Configuration: $M$ = DinoV2-G on AGX Orin.
As expected, pruning earlier yields lower latency but greater accuracy degradation. For the batch size 1, our method pruned $\sim$54of input tokens at the first layer degraded accuracy by 3.19% but yielded a 40% overall latency reduction. Across both batch size 1 and 2 in this ablation study, pruning after the first 25% of layers (layer 10) results in a good balance between latency reduction and accuracy degradation. Pruning later in the network will reduce accuracy degradation, however we prioritize yielding latency benefits with our method. Therefore, we perform pruning at the layer 25% of the way into the network for all models evaluated in this work. That is layer 3 for DeiT-B and DeiT-S, and layer 10 for DinoV2-G.
Device Batch Model & Acc. Median
Size Method Loss Latency (ms)
@@ 4.3 Pruning location ablation study
TX2 2 DeiT-S -(00.0%) 68.92
w/ Top-K 4.30 (-27.0%) 50.32
w/ ToMe 2.47 (-23.1%) 52.98
w/ DyViT ✝ 0.46 (-26.2%) 50.88
w/ Ours 1.24 (-28.3%) 49.44
@@
Table 6: Comparison with existing methods for similar latency across device and workload sizes. We tune the pruning hyperparameters of the compared methods so they achieve similar latency to our method. ✝DynamicViT (DyViT) requires training. Other evaluated methods are training-free.
4.4 Accuracy-latency tradeoffs of our method
In this section, we discuss the results of our training-free token pruning mechanism when combined with the pruning schedule inspired by the GPU tail effect. We primarily compare against ToMe [1], and include Top-K as a baseline [13, 19]. DynamicViT [36] is included in select evaluations, however unlike the other compared work and our method it requires training. Tab. 6 illustrates our method’s ability to retain higher accuracy for similar latency across various devices and workload sizes. Fig. 6 demonstrates that our token pruning mechanism and schedule expands the accuracy-latency tradeoffs on the pareto front across various workload sizes and devices.
First, our method is able to achieve higher accuracy than other training-free pruning techniques ToMe and Top-K. Tab. 6 is an ablation study across various workload sizes (batch size, models) and devices, where we tune the hyperparameters of methods such that similar latency is achieved. Unsurprisingly, DynamicViT is able to retain higher accuracy than all methods, since it requires training. Compared with ToMe and Top-K however, it is shown that our method yields consistently lower accuracy degradation. Batch size 2 inference on the AGX Orin demonstrates that to achieve similar latency to our method, Top-K degrades accuracy by more than 45.6%, while we degrade accuracy by only 2%. Across all workload sizes and devices, ToMe featured 0.47%-5.92% lower accuracy than our method (Tab. 6).
@@ 4.4 Accuracy-latency tradeoffs of our method
Second, we note that for larger pruning rates such as $r=7$ and $r=8$, ToMe and Top-K remove nearly all input tokens by the last layer of DinoV2-G. These high pruning ratios yield high accuracy degradation of over 40% in the case of A100 batch-size 4 for ToMe, as seen in Fig. 6. Comparatively, by pruning 54%-75% of tokens at the 10th layer of DinoV2-G according to the tail effect, we achieve higher accuracy and lower latency than progressive pruning approaches. This shows aggressive pruning at early layers can be strictly better than progressive pruning methods for certain workload sizes.
4.5 Low workload size observations
We demonstrate there are some cases where pruning techniques actually increase latency with respect to the baseline model, rather than reduce it. Tab. 7 demonstrates an example on the TX2 for single batch inference. ToMe increased latency by $\sim$30%, while we reduced latency by $\sim$9%. Our tradeoff of 0.99% accuracy degradation and 9% latency reduction is better than ToMe and Top-K, both of which degrade accuracy and provide no latency benefit compared with the baseline DeiT-S pre-trained model. For certain workload sizes, using a baseline model or a hardware aware method such as ours is strictly better than attempting to use a token removal method with high overhead.
Device Method Acc. Median
Loss Latency (ms)
TX2 Baseline (-0.00%) 35.32
Top-K (r=6) 0.73 (+18.6%) 43.38
ToMe (r=6) 0.27 (+30.3%) 50.70
Ours (R=69) 0.99 (-9.06%) 32.12
Table 7: Experiment on DeiT-S with batch size 1 that illustrates pruning overhead. Given comparable total pruned tokens (69 to 72), for certain workload sizes a token pruning mechanism may increase latency due to the overhead of performing pruning at multiple layers. Our pruning method is applied at a single layer to avoid this relatively significant overhead.
We found that the operations required to do token pruning add a non-trivial latency overhead for small workload sizes on some devices such as the TX2. The overhead from masking or removing tokens can outweigh the latency reduction from processing fewer tokens. As explained in Sec. 3.2, our choice to perform single-layer pruning rather than progressive pruning allows us to yield a latency reduction.
@@ 5 Limitations
Our approach to token pruning is designed to achieve better tradeoffs when latency-workload is non-linear. This occurs primarily for small workload sizes and batch sizes, which are common contexts in edge deployment. For large workloads such as DinoV2-G with batch size 256+, our method is less effective because latency-workload relationships becomes linear. Last, though we avoid training, our method requires a grid-search. In some contexts, this may be expensive (many tokens, large batch size, large model). These contexts are outside the focus of this work, but may be relevant for some practitioners.
@@ 6 Conclusion
In this work, we investigated how to improve the efficiency of vision transformers by considering latency-workload relationships. We made three main contributions. First, we provide insight on how ViT latency is affected by underlying hardware, machine learning framework, and kernel launch factors. Second, we show how to determine a token pruning schedule by leveraging information about latency-workload relationships. Third, we introduced a token pruning mechanism that is training-free and low overhead. Our method advances the accuracy-latency Pareto front compared with existing token pruning approaches across workload sizes and devices. For single batch inference, other methods may increase latency by 18.6-30.3% with respect to baseline, while we reduce it by 9%. For similar latency (within 5.2%) across devices, the state-of-the-art, ToMe, achieves 45.8%-85.4% ImageNet1K accuracy, while we achieve 78.6%-84.5%. Improving kernel designs to reduce non-linear latency behavior is a promising future direction.

### v3

@@ 3.3 Token Pruning Method
The schedule selection described in Secs. 3.1 and 3.2 identifies the location and number of tokens to prune. Now, we give a training-free token pruning mechanism to decide which tokens to prune at inference as in Fig. 3. The offline pruning schedule, consisting of a number of tokens to prune $R$ and the layer index $L$ at which to prune, is an input to our token pruning mechanism. The primary design goal for our pruning mechanism is to require no finetuning on $M$ and be lightweight - thus we restrict ourselves to using intermediate computations from the attention operation.
First, we choose a method to rank the importance of tokens. Following prior work [2, 14], we rank token importance by measuring the attention each token receives from all others, utilizing the softmax attention matrix. We also incorporate an importance term derived from the $V$ matrix, which marginally increased accuracy.
Second, we borrow from EViT [18] and instead of discarding pruned tokens, we create a new “inattentive” token based on the features of all pruned tokens, then and append it to the set of kept tokens. Information is thus preserved from the pruned tokens, increasing accuracy while reducing the total token count.
Algorithm 3 depicts a forward pass using our pruning mechanism. Inputs from our offline computation, $R$ and $L$, are utilized to rank tokens and decide at which layer to prune. rank_tokens ranks each token based on our V matrix importance vm and a standard attention matrix importance term am. prune_tokens prunes $R$ tokens using a standard token removal implementation [34].
3.4 Qualitative Comparison with Pruning and Merging
Here, we justify why we classify our method as token pruning rather than merging. Our method uses the “inattentive“ token from EViT, which the EViT authors consider a hybrid method. However, we take the position that the core mechanism of deciding which tokens to prune is an important differentiating factor. A majority of pruning-based approaches treat the selection of tokens to prune as a ranking problem, rather than a matching problem as merging methods do. Our importance score computation is most similar to ranking-based approaches. Thus, we see ourselves primarily as a pruning method, though we could be considered a hybrid between pruning and merging.
@@ 4 Evaluation
After describing experimental setup (Sec. 4.1), we characterize our technique via ablation over $\alpha$ and by measuring offline computation costs (Sec. 4.2). Then we compare to the state-of-art ToMe method and relevant baselines (Sec. 4.3).
4.1 Experimental Setup
Hardware: We use three devices with varying characteristics (Tab. 3). Our technique targets edge workloads, so we use two edge-caliber development boards designed for machine learning: NVIDIA TX2 [28] and NVIDIA AGX Orin [23]. To assess generalization beyond edge devices, we also use a server-grade NVIDIA A100 GPU [24]. On the TX2 and Orin we used fixed CPU and GPU clock rates for consistency. The A100 system clocks could not be locked because those servers are shared resources.
@@ 4.1 Experimental Setup
Model Params (M) Depth
DeiT-S [37] 21 12
DeiT-B [37] 86 12
ViT-L [9, 42] 300 24
DinoV2-G [29] 1100 40
Table 4: Characteristics of models we evaluate over. We include DinoV2 as a representative state-of-the-art ViT, while ViT and DeiT are commonly used baselines in prior token pruning work.
Models: Tab. 4 summarizes the models used. We evaluate common vision transformer models across a variety of scales (21M to 1.1B parameters). For DeiT and ViT models we use the TIMM pretrained weights, while we use the DinoV2-G weights from the DinoV2 Github.
Measurements: In order to measure latency for evaluation, we use the PyTorch benchmark module [32]. Latency is measured over 16 seconds. To be clear, we define latency as the compute time required for a forward pass of a model given a batch of input images. Accuracy was measured using the A100 system, with a batch size of 512 on the evaluation subset of ImageNet1K. In subsequent experiments, we decide the pruning hyperparameters of each token pruning method for fair comparison with our method. For reproducibility, our code is open source.
4.2 Characterizations of our technique
Here we evaluate two of the three design decisions of our method. First, we ablate over the hyperparameter $\alpha$ that is used to weight the utility function. Second, we measure the cost (time) for offline computation. The third decision, selecting the layer $L$ for pruning, is evaluated in supplemental material.
Pruning schedule ablation study across $\alpha$. In Sec. 3.1 we introduce an algorithm to decide a number of tokens to prune $R$ according to the GPU tail effect.
[…]
@@ 4.2 Characterizations of our technique
The $\alpha$ hyperparameter governs the relative weighting of accuracy and latency utilities in our algorithm. Setting $\alpha=1$ prioritizes accuracy, and only a few tokens might be pruned. Setting $\alpha=0$ prioritizes latency, and would prune all or nearly-all tokens. To decide the value of $\alpha$, we performed an ablation study in  Tab. 5.
When evaluated on the AGX Orin with $\alpha>0.5$, we computed $R=193$ (over 75% of tokens pruned), resulting in a $\sim$2.8accuracy drop for a 2.1% latency reduction. We consider this to be an unfavorable tradeoff. This ablation, in addition to results in Sec. 4.3, suggest that $\alpha\leq 0.5$ is a good choice. Thus we use $\alpha=0.5$ for all evaluations that appear in this work. Intuitively speaking this means accuracy degradation and latency reduction are considered with equal weight according to Eq. 3 when selecting $R$.
Offline computation time. In  Sec. 3.1, we describe our offline computations to decide a number of tokens to prune in which we utilize a grid search to measure latency and estimate accuracy degradation.  Tab. 6 illustrates the total times required for measuring the workload-latency characteristics of DeiT-S and ViT-L across devices. The offline computation is relatively fast (no more than 4.5 hours), especially compared with the time required to train any ViT.
[…]
Time (min) Time (min)
TX2 DeiT-S - 54
ViT-L - 54
Orin DeiT-S - 54
ViT-L - 54
A100 DeiT-S 83 54
ViT-L 218 52
[…]
Size Loss Latency (ms)
TX2 2 DeiT-S -(00.0%) 68.92
w/ Top-K 4.30 (-27.0%) 50.32
w/ ToMe 2.47 (-23.1%) 52.98
w/ DyViT ✝ 0.46 (-26.2%) 50.88
w/ Ours 1.24 (-28.3%) 49.44
@@
TX2 2 ViT-L -(00.0%) 1327.0
w/ Top-K 38.2 (-57.2%) 568.0
w/ ToMe 17.5 (-57.8%) 559.3
w/ Ours 8.4 (-55.2%) 594.9
Orin 4 ViT-L -(00.0%) 70.29
w/ Top-K 52.70 (-32.2%) 47.63
w/ ToMe 17.51 (-22.9%) 54.18
w/ Ours 2.35 (-33.0%) 47.08
@@
Table 7: Comparison with existing methods for similar latency across devices and workload sizes. Hyperparameters of methods were chosen to get as close as possible to match our latency; in some cases pruning more tokens did not reduce latency. For each model, the smallest batch size was selected to demonstrate cases where the workload-latency relationship can be exploited. ✝DynamicViT (DyViT) is the the only training-based method.
@@
4.3 Comparison to Other Methods
In this section, we demonstrate the effectiveness of our token pruning schedule and pruning mechanism across devices and workload sizes. Our primary focus is on smaller workloads, which are typical in edge deployment scenarios [17].
For inter-method comparison, we systematically compare to the state-of-the-art method, Token Merging (ToMe) [1]. We also evaluate two common benchmarks, Top-K [12, 18] and DynamicViT [34]. We measure Top-K in all conditions; we use DynamicViT only with models for which its pre-trained weights were available. DynamicViT is not emphasized as it involves training, making it an unfair comparison with our method, ToMe, and Top-K.
Accuracy-latency tradeoffs. In this section, we discuss the results of our training-free token pruning mechanism and offline computed pruning schedule. Tab. 7 illustrates our method’s ability to retain higher accuracy for similar latency across devices and workload sizes. Fig. 4 demonstrates that our token pruning mechanism and schedule expands the accuracy-latency tradeoffs on the pareto front across devices and workload sizes.
First, our method is able to achieve higher accuracy than other training-free pruning techniques ToMe and Top-K. Tab. 7 is an ablation study across various workload sizes (batch size, models) and devices, where we tune the hyperparameters of methods such that similar latency is achieved. Unsurprisingly, DynamicViT retains accuracy since it was finetuned for 300 epochs. However, compared to ToMe and Top-K, our method consistently results in lower accuracy degradation. In one case, Top-K degrades accuracy by more than 45.6%, while we degrade accuracy by only 2%. Across all workload sizes and devices, ToMe had 0.47 to 15.16 lower Top-1 percentage points than our method for similar latency (within 5.2% or 7ms).
@@ 4.3 Comparison to Other Methods
Second, we note that for larger pruning rates such as $r=7$ and $r=8$, ToMe and Top-K remove nearly all input tokens by the last layer of DinoV2-G. These high pruning ratios yield high accuracy degradation of over 40% in the case of A100 batch-size 4 for ToMe, as seen in Fig. 4. Comparatively, by pruning 54%-75% of tokens at the 10th layer of DinoV2-G according to the tail effect, we achieve higher accuracy and lower latency.
In both Tab. 7 and Fig. 4, our method features lower accuracy degradation than ToMe and Top-K. At small workloads, the marginal latency of pruning additional tokens becomes negligible. As a result, pruning tokens at each layer degrades accuracy significantly for little latency benefit; existing methods do not account for this behavior. Thus, we prune $R$ tokens early; the remaining tokens propagate through the ViT, retaining information which leads to better accuracy. Simultaneously, we achieve high latency reduction due to pruning early in the network.
Low workload size observations. For small workloads, token sparsification can actually increase latency due to the overhead associated with token removal mechanisms. Tab. 8 illustrates three examples of this. ToMe and Top-K may increase latency by 2-30% with respect to baseline, while we reduce it by 9-26%. There are also cases where all methods, including ours, increase latency by 40%-134%. We find that for some workloads, using a baseline model or our method is strictly better than attempting to use a token removal method with high overhead.
[…]
Size Loss Latency (ms)
TX2 1 DeiT-S (-0.00%) 35.32
Top-K 0.73 (+18.6%) 43.38
ToMe 0.27 (+30.3%) 50.70
Ours 0.99 (-9.06%) 32.12
A100 4 ViT-L (-0.00%) 9.49
Top-K 0.47 (+51.4%) 14.36
ToMe 0.29 (+134.4%) 22.24
Ours 0.85 (+40.0%) 13.29
A100 16 ViT-L -(00.0%) 29.41
w/ Top-K 0.77 (+2.48%) 30.14
w/ ToMe 0.51 (+7.65%) 31.66
w/ Ours 2.26 (-26.3%) 21.68
Table 8: Experiment that illustrate pruning overhead for certain workload sizes. Pruning parameters were chosen such that a similar number of tokens were pruned as our method. Token pruning methods may increase latency due to the overhead of pruning itself. We reduce overhead through single-layer pruning.
@@ 5 Limitations
Our token pruning approach is optimized for cases where latency-workload relationships are non-linear. For large workloads such as DinoV2-G with batch size 256+, our method is less effective because latency-workload relationships becomes linear and more predictable in this case.
@@ 6 Conclusion
In this work, we offered practical guidance on how to improve token pruning for ViTs in the presence of small workloads by utilizing latency-workload relationships. We showed how to determine a token pruning schedule by leveraging non-linear latency-workload relationships; in comparison with prior work, our method yields equal or greater latency reductions while maintaining greater accuracy. Ultimately, we demonstrated that leveraging workload-latency behavior is effective at improving ViT efficiency via token pruning, especially for small workloads.